### v2

# Mechanism Shift During Post-training from Autoregressive to Masked Diffusion Language Models

Abstract
Post-training pretrained Autoregressive models (ARMs) into Masked Diffusion models (MDMs) has emerged as a cost-effective strategy to overcome the limitations of sequential generation. However, the internal algorithmic transformations induced by this paradigm shift remain unexplored, leaving it unclear whether post-trained MDMs acquire genuine bidirectional reasoning capabilities or merely repackage autoregressive heuristics. In this work, we address this question by conducting a comparative circuit analysis of ARMs and their MDM counterparts. Our analysis reveals a systematic “mechanism shift” dependent on the structural nature of the task. Structurally, we observe a distinct divergence: while MDMs largely retain autoregressive circuitry for tasks dominated by local causal dependencies, they abandon initialized pathways for global planning tasks, exhibiting distinct rewiring characterized by increased early-layer processing. Semantically, we identify a transition from sharp, localized specialization in ARMs to distributed integration in MDMs. Through these findings, we conclude that diffusion post-training does not merely adapt model parameters but fundamentally reorganizes internal computation to support non-sequential global planning.
Mechanism Shift During Post-training from Autoregressive to Masked Diffusion Language Models
Injin Kong1,∗ Hyoungjoon Lee2,∗ Yohan Jo1,† 1Graduate School of Data Science, Seoul National University 2Department of Biosystems & Biomaterials Science and Engineering, Seoul National University mtkong77@snu.ac.kr, hjoon721@snu.ac.kr, yohan.jo@snu.ac.kr
1 Introduction
Large language models have achieved near-human performance across diverse linguistic tasks OpenAI et al. (2024); Qwen et al. (2025); Touvron et al. (2023). Despite these advances, the prevalent autoregressive framework Vaswani et al. (2017); Radford et al. (2019) imposes structural limitations on generation Welleck et al. (2019); Bengio et al. (2015). Specifically, sequential generation under causal masking prevents correcting past tokens Gu et al. (2018); Welleck et al. (2019); Vaswani et al. (2017) so that early errors propagate and amplify throughout the sequence, as the model cannot correct past inaccuracies Bengio et al. (2015); Ranzato et al. (2016). Moreover, many reasoning and planning tasks require global reasoning, where early decisions must account for constraints that apply to the entire sequence Gu et al. (2018); Ye et al. (2025a).
Masked diffusion models (MDMs) have gained increasing interest as a non-autoregressive paradigm, with structural properties well-suited to overcoming these limitations Austin et al. (2021); Sahoo et al. (2024). However, training diffusion-based language models from scratch remains computationally expensive due to slower convergence Gong et al. (2025). To mitigate this cost, recent work proposes post-training pretrained autoregressive models (ARMs) to the diffusion paradigm Gong et al. (2025); Ye et al. (2025b). Models such as Dream Ye et al. (2025b) demonstrate that this strategy can achieve strong performance while requiring only a fraction of the compute needed for training from scratch.
Despite the empirical success of post-training ARMs with diffusion objectives, the specific algorithm changes induced by this post-training process are not yet understood Gong et al. (2025). It remains unclear whether post-trained MDMs genuinely learn new bidirectional reasoning mechanisms, as intended by the diffusion framework, or instead still rely heavily on autoregressive mechanisms at their core. Although mechanistic interpretability has been applied to understand diffusion models for image generation Shi et al. (2025); Niedoba et al. (2025), this level of analysis has not yet been extended to text diffusion. Without such analysis, it is difficult to determine if post-trained MDMs genuinely perform global reasoning. If models continue to rely on local, left-to-right heuristics, the theoretical benefits of diffusion-based architectures fail to materialize.
In this work, we address this question by analyzing language models from a circuit-level perspective Bhaskar et al. (2024), which allows us to directly examine whether diffusion post-training induces new computational pathways or primarily reuses existing autoregressive ones. We investigate where algorithmic changes occur by comparing circuit structures between ARMs and MDMs post-trained from the same autoregressive backbones, and then examine how these changes are realized through detailed analysis using logit lens techniques and neuron-level visualizations.
Through this analysis, we demonstrate that post-training using MDM objectives does not merely alter the training loss but instead induces a systematic reorganization of internal computation—shifting semantic roles across components and revealing a mechanism shift in how language models process and refine linguistic information.
2 Related Works
2.1 Masked Diffusion Models
MDMs generate text by reversing a corruption process that stochastically replaces tokens with a [MASK] symbol Chang et al. (2022); Austin et al. (2021). This approach allows for non-autoregressive generation using full bidirectional context.
While training such models from scratch Nie et al. (2025) is computationally expensive, recent work has shown that pretrained ARMs can be effectively post-trained into MDMs Gong et al. (2025). Rather than learning diffusion dynamics from scratch, these approaches initialize from a pretrained ARM. The model is then post-trained to iteratively denoise partially masked inputs, instead of predicting the next token autoregressively. From DiffuLLaMA Gong et al. (2025) to Dream Ye et al. (2025b), this line of work shows that post-training pretrained ARMs to MDM objectives can retain many practical advantages of diffusion—such as parallel decoding, iterative refinement, and bidirectional attention—while substantially reducing training cost. Post-trained MDMs also achieve strong performance on directionality-sensitive tasks. However, while these works establish the effectiveness of post-trained MDMs, they largely focus on performance and efficiency, leaving open the question of how diffusion objectives reshape the underlying computational mechanisms.
2.2 Mechanistic Interpretability and Circuits
Mechanistic interpretability aims to identify the internal components and algorithms responsible for specific model behaviors Olah et al. (2020); Elhage et al. (2021). A central concept is the circuit, defined as a subgraph of the computational graph connecting inputs to the unembedding projection that is sufficient to produce a target behavior Olah et al. (2020); Bhaskar et al. (2024). Nodes correspond to components such as attention heads and MLPs, while directed edges represent causal dependencies between components, where the output of one node contributes to the input of another Ou et al. (2025); Hanna et al. (2024).
Empirical studies show that many behaviors can be explained by sparse circuits involving only a small fraction of model connections Bhaskar et al. (2024); Wang et al. (2023). Methods such as Edge Attribution Patching identify these subgraphs via gradient-based attribution, enabling circuit discovery for tasks including indirect object identification and numerical comparison Hanna et al. (2023); Lieberum et al. (2023); Bhaskar et al. (2024). Across models and scales, similar circuits—such as induction heads—recur consistently, suggesting that they implement stable algorithmic functions rather than incidental patterns Prakash et al. (2024); Tigges et al. (2024); Ou et al. (2025); Wang et al. (2025). Recent automated approaches, including ACDC and EAP, further enable scalable circuit discovery without manual inspection Syed et al. (2024); Bhaskar et al. (2024).
While mechanistic analyses have begun to probe diffusion models in the vision domain Shi et al. (2025); Niedoba et al. (2025), comparable studies for text diffusion models remain limited. As a result, it is unclear whether diffusion objectives induce distinct computational strategies in language models or primarily reorganize existing autoregressive circuitry.
3 Method
To investigate the mechanistic shift from ARMs to MDMs, we adopt a comparative framework. Our primary objective is to distinguish the impact of shifting the learning objective from causal modeling to masked diffusion on circuit topology. We control for architectural confounding by analyzing ARMs alongside their directly post-trained MDM counterparts.
3.1 Models and Configuration
We conduct experiments across two distinct model families to verify the generalizability of our findings. Specifically, we utilize the Qwen2.5-7B Qwen et al. (2025) and LLaMA-2-7B Touvron et al. (2023) architectures.
To enable a direct comparison, we pair each standard autoregressive checkpoint with an MDM post-trained from that specific ARM base:
3.2 Tasks and Datasets
We analyze circuit behavior on two distinct tasks to highlight different performance capabilities: one focused on autoregressive causal dependencies and another on global reasoning. Additional details for tasks and datasets are provided in Appendix A.
Indirect Object Identification (IOI):
A canonical interpretability task primarily solved by Induction Heads in ARMs Wang et al. (2023). We select this task to serve as a representative baseline for causal reasoning; specifically, we aim to investigate whether the induction circuitry—which inherently relies on sequential, left-to-right context to copy tokens—persists, dissolves, or transforms when the training objective shifts from causal prediction to masked diffusion.
Countdown:
A numerical reasoning task where the model is provided with a set of integers and a target value and must generate a valid arithmetic equation using the inputs to equal the target Ye et al. (2025a). This task serves as a critical benchmark for global reasoning because it requires inverse planning—decisions made early in the sequence must be strictly conditioned on the final goal—thereby challenging the left-to-right causality of ARMs while favoring the bidirectional context and global refinement capabilities of MDMs.
Inference Configuration
For all experiments, we use task-dependent generation lengths. For IOI, we use a single diffusion step, while for Countdown, we align diffusion steps with the target sequence length. Details are provided in Appendix A.
3.3 Circuit Discovery and Analysis Pipeline
We hypothesize that the transition to masked diffusion induces specific algorithmic shifts in tasks where MDMs outperform ARMs (e.g., Countdown). To verify this, we employ a three-stage pipeline: Discovery, Topological Comparison, and Mechanism Interpretation. This workflow allows us to first locate where the computation changes and then analyze what those changes represent semantically.
3.3.1 Discovery
We employ automated circuit discovery based on Edge Attribution Patching with Integrated Gradients(EAP-IG) Hanna et al. (2024) to identify the minimal computational subgraph responsible for task performance. For a given task, we identify a sparse subgraph containing the subset of edges required to keep the model’s performance within a threshold of the full model. Within this discovered circuit, nodes represent distinct functional components: attention heads primarily act as information movers that copy content from preceding tokens, while MLPs often serve as associative memories that extract or refine specific semantic attributes from the input state. Consequently, the edges structurally define the algorithmic logic, specifying how these functional outputs are composed and routed to produce the final prediction.
3.3.2 Attribution-Guided Circuit Comparison
To localize where algorithmic changes concentrate when transitioning from ARMs to MDMs, we leverage EAP-IG to compare circuits at the level of edges and their incident components. Rather than treating all discovered edges equally, we focus on those that carry the highest attribution mass under EAP-IG and then identify which components are repeatedly used as sources or sinks of these high-attribution edges.
EAP-IG Edge Overlap:
For each model (ARM and MDM), we first select the set of top-attribution edges (1000 edges), and , by ranking edges according to their EAP-IG scores on identical prompts. We then quantify overlap using the Jaccard similarity . High overlap suggests that the MDM reuses the same high-attribution pathways as the ARM, whereas low overlap indicates that diffusion training recruits a distinct set of edges to implement its generative computation. The choice of selecting the top 1000 edges reflects a trade-off between attribution coverage and circuit sparsity and is empirically justified in Appendix B.
Top-K EAP-IG Components:
To move from edges to components, we assign each node a score by aggregating the EAP-IG scores of all incident edges, defined as the sum of incoming attributions and outgoing attributions , where denotes the set of top-attribution edges for the model. In words, a component is important if it repeatedly appears as either the source or the target of high-scoring EAP-IG edges. We then define the Top-K Components (with ) for each model as the nodes with the largest . The choice of is empirically motivated and discussed in Appendix B.
Based on these metrics, we first assess the degree of circuit reuse between ARMs and MDMs via overlap of high-attribution edges. High edge overlap provides evidence of mechanistic recycling, where the MDM relies on pre-trained autoregressive pathways, while low overlap indicates the emergence of diffusion-specific generative strategies. Within these regions, the Top-K EAP-IG Components pinpoint nodes that sit at the endpoints of the most influential edges, yielding a focused set of components for downstream mechanistic analysis via logit lens and neuron-level visualization.
3.3.3 Mechanism Interpretation
Having identified where the circuit changes, we investigate how the computation differs by applying interpretability techniques specifically to the Top-K Divergent Components:
Logit Lens and Component-wise Analysis:
Following the logit lens framework nostalgebraist (2020), we project intermediate activations into the vocabulary space via the unembedding matrix , yielding and , where indexes a component (e.g., an attention head), denotes the token position, and the diffusion timestep.
For autoregressive models (Qwen, LLaMA), we apply the component-wise logit lens to all token positions in the sequence on both IOI and Countdown tasks. This enables us to trace how the model’s internal representation becomes aligned with the target output across the sequence and to decompose the contributions of the residual stream, attention heads, and MLPs at each token position.
For MDMs (Dream, DiffuLLaMA), we examine how these component-wise projections evolve across the diffusion time steps at a fixed set of token positions. Dream supports per-head and per-MLP decompositions, while DiffuLLaMA is analyzed at the level of residual, attention, MLP, and residual-out layers.
This analysis allows us to assess whether a given component has shifted its semantic role—for example, from contributing primarily to next-token prediction in an ARM to encoding a global target or distant operator in an MDM—and how such semantic alignment emerges progressively over diffusion time. Importantly, when applying the unembedding matrix to intermediate component activations, we do not interpret the resulting logits as the model’s final predictions. Instead, they serve as a diagnostic probe that reveals which tokens a component is linearly aligned with at a particular stage of computation.
Neuron Explanation:
In this work, we define a “neuron” as a single scalar coordinate in the model’s residual stream at a given transformer layer (i.e., one element of the hidden dimension of the layer output). For both the IOI and Countdown settings, we record activations for all combinations of layers, neuron indices, and token positions on the evaluation data. Following the methodology of Bills et al. (2023), we then extract the input tokens that elicit the largest-magnitude activations for a given neuron and use these highly activating examples for qualitative inspection and automated explanation.
We focus this fine-grained analysis on the Top-K Divergent Components identified in the previous step: specifically, for divergent MLP layers, we visualize the top activating neurons to understand the specific attributes they extract; for divergent Attention layers, we analyze the head’s output features to determine what information is being moved. The primary goal of this visualization is to characterize the semantic feature selectivity of the circuit. By mapping high-activation neurons to their corresponding tokens, we aim to determine if MDM components have learned to encode non-causal features (e.g., attending to or encoding “future target” tokens available during the diffusion process) that are fundamentally inaccessible to the standard autoregressive model.
4 Results & Analysis
4.1 Circuit-Level Differences
| Setting | Edge Overlap | Top- Overlap |
| IOI (Qwen / Dream) | 0.193 | 0.105 |
| IOI (LLaMA-2 / DiffuLLaMA) | 0.088 | 0.124 |
| Countdown (Qwen / Dream) | 0.008 | 0.093 |
| Countdown (LLaMA-2 / DiffuLLaMA) | 0.032 | 0.081 |
Figure 2 illustrates the circuit structures extracted for both ARMs and MDMs across the IOI and Countdown tasks, where each circuit represents the average structure aggregated over multiple prompts and diffusion steps.
For the IOI task, the overall circuit topology of Dream closely resembles that of Qwen, with dominant interactions concentrated in similar layers and mediated by comparable attention pathways. A similar pattern is observed for DiffuLLaMA relative to its LLaMA-2 counterpart. This indicates that, for tasks primarily driven by local or causal dependencies—such as IOI, which effectively requires only shallow or single-step generation—post-trained MDMs largely preserve the autoregressive circuitry inherited from their ARM initializations.
This qualitative similarity is further supported by quantitative circuit similarity metrics reported in Table 1. Across both model families, the IOI task exhibits consistently higher overlap in high-attribution edges and Top- components compared to the Countdown task, indicating substantial mechanistic reuse rather than the emergence of new diffusion-specific circuitry in tasks dominated by local dependency structures.
In contrast, the Countdown task exhibits a clear shift in circuit organization between ARMs and MDMs. In the Countdown task, the ARM circuit becomes increasingly dense toward mid-to-late layers, whereas the MDM circuit concentrates a larger fraction of interactions in early layers, resulting in a front-loaded organization of computation. This redistribution is consistently observed both in the aggregated circuit statistics (Figure 2) and in the representative circuit visualizations (Figure 3), indicating a systematic reallocation of computation under the diffusion objective.
This depth-wise shift toward increased early-layer engagement is consistently observed across both model families, with Dream and DiffuLLaMA exhibiting increased early-layer engagement compared to their respective autoregressive baselines. Quantitatively, this divergence is reflected in markedly lower edge-level and component-level overlap scores for the Countdown task (Table 1), indicating that diffusion post-training induces non-trivial reorganization of the underlying circuitry.
Together, these results reveal a clear dichotomy in how masked diffusion post-training reshapes internal circuitry. For tasks dominated by local, causal dependencies such as IOI, post-trained MDMs retain a substantial fraction of the original autoregressive circuitry. High overlap in both edge-level attribution and Top- components indicates that the causal reasoning machinery learned during autoregressive pre-training remains functionally effective and is therefore reused rather than replaced.
In contrast, for tasks requiring global constraint satisfaction, such as Countdown, the autoregressive circuitry proves insufficient. Here, diffusion post-training induces the emergence of distinct early-layer components that are largely absent from the original ARM circuits, accompanied by a marked collapse in circuit overlap. These newly emphasized early-layer components become the primary substrates for computation, suggesting that MDMs restructure their internal mechanisms to support global planning by front-loading computation rather than relying on sequential, causal pathways.
Taken together, these findings demonstrate that masked diffusion post-training does not simply overwrite autoregressive mechanisms. Instead, it selectively preserves causal circuitry where it remains compatible with the task, while inducing genuinely new, diffusion-specific components when global reasoning demands exceed the expressive capacity of autoregressive computation.
4.2 Semantic Reorganization of Task-Critical Components and Early Layers
| Task | Model | Component | Mean Logit | Top Tokens | Role |
| IOI | Qwen | a27.h5 | 21.14 | Dan | Person-name-related Component |
| Qwen | a23.h11 | 6.84 | Ben | Person-name-related Component | |
| LLaMA | a24.h15 | 2.75 | Jerry | Person-name-related Component | |
| LLaMA | a21.h1 | 1.73 | Carol | Person-name-related Component | |
| Dream | m25 | 3.14 | Browser | Proper-noun component | |
| Dream | m22 | 2.33 | Kremlin | Proper-noun component | |
| DiffuLLaMA | a26.h21 | 2.81 | Marian | Person-name-related Component | |
| DiffuLLaMA | a22.h19 | 1.80 | Grace | Person-name-related Component | |
| Countdown | Qwen | m25 | 51.29 | 3 | Digit-related Component |
| Qwen | m20 | 20.67 | 1, 2 | Broad Numerical Component | |
| LLaMA | m29 | 6.20 | pick | Instruction-related Component | |
| LLaMA | a22.h13 | 2.10 | four | Numerical–Lexical Component | |
| Dream | m27 | 26.06 | 1, 2, 3 | Broad Numerical Component | |
| Dream | m23 | 5.34 | 5, 1, 4 | Broad Numerical Component | |
| DiffuLLaMA | m31 | 12.32 | 0, 1, 2 | Broad Numerical Component | |
| DiffuLLaMA | m4 | 1.40 | – | Symbol-related Component |
To understand how circuit-level differences translate into concrete computational strategies, we analyze model behavior at two complementary granularities. First, we examine task-critical components—attention heads and MLPs that receive high EAP-IG attribution—using component-wise logit lens analysis. This allows us to characterize how semantic roles are assigned to components that directly influence task outputs. Second, we analyze early-layer neuron activations to understand how these semantic roles are implemented internally, particularly in regions where MDMs emphasize computation but component-level semantics appear diffuse.
4.2.1 Task-Critical Components: Component-wise Logit Lens Analysis
We begin by analyzing components that directly contribute to task performance, as identified by high EAP-IG attribution. By applying a component-wise logit lens to these components, we probe their semantic alignment with output tokens, allowing us to assess whether task-relevant information is concentrated within a small number of specialized components or distributed across many.
Causal reasoning task (IOI)
For the IOI task, ARMs exhibit sharply localized semantic specialization. In both Qwen and LLaMA, a small number of attention heads are strongly aligned with person-name tokens (e.g., Dan, Jerry), producing large-magnitude logits that dominate prediction. This pattern indicates that IOI is largely solved by a limited set of highly specialized components that act as deterministic, pointer-like mechanisms.
DiffuLLaMA largely preserves this behavior. Several high-attribution components remain strongly aligned with person-name tokens and exhibit logit profiles comparable to those of the underlying ARM. This suggests substantial inheritance of autoregressive circuitry, consistent with the observation that IOI relies primarily on local causal dependencies that are already well supported by the autoregressive objective.
Dream, however, departs from this pattern. Its high-attribution components do not exhibit clear person-name specificity. Instead, the strongest logit alignments correspond to non-person proper nouns or broader semantic tokens (e.g., Browser, Kremlin). This indicates that, despite similar circuit topology at a coarse level, diffusion post-training redistributes semantic roles across components, weakening the dominance of individual name-specific heads.
Global reasoning task (Countdown).
A contrasting pattern emerges in the Countdown task. ARMs rely on components with strong numerical selectivity, often sharply aligned with specific digits or operators. These components produce high-confidence logits concentrated on a narrow subset of tokens, indicating that global planning is approximated through sequential, component-centric heuristics.
In Dream, this sharp specialization collapses. Instead of a single dominant numerical component, multiple components exhibit moderate logit responses distributed across numerical tokens, with no individual component exerting decisive control over prediction. DiffuLLaMA again occupies an intermediate regime, retaining numerical associations but with reduced magnitude and increased dispersion.
Taken together, these results indicate that diffusion post-training induces a semantic reorganization at the component level. Whereas ARMs resolve tasks through a small number of highly specialized components, MDMs—particularly on global reasoning tasks—redistribute semantic responsibility across multiple components, yielding a more ensemble-like computation.
4.2.2 Early-layer Representation: Neuron-level Analysis
While component-wise logit lens analysis captures semantic specialization among task-critical components, it does not fully explain the behavior of early-layer regions that are emphasized by masked diffusion circuits. To characterize how task-relevant information is implemented in these layers, we analyze neuron activations in the lowest transformer layers across tasks and model classes.
In ARMs, neuron activation in early layers is highly task-dependent. In the IOI task, a large number of neurons in early layers are strongly activated by descriptive adjectives and modifier-related tokens, suggesting that early layers encode fine-grained syntactic and descriptive features prior to resolving entity identity. In the Countdown task, this pattern shifts markedly toward neurons associated with numerical or technical content. Despite this task-dependent reallocation, ARMs consistently exhibit strong concentration within specific semantic categories.
MDMs display a qualitatively different pattern. Across both IOI and Countdown, the total number of active neurons in early layers remains relatively stable. Moreover, the activated neurons tend to correspond to broad, genre-level cues—such as general numerical or technical content—rather than sharply defined task-specific categories. This indicates that MDMs rely less on early-layer specialization and instead maintain a more uniform, task-agnostic representational regime.
This contrast persists in the early layer. Across depth, ARMs repeatedly exhibit sharper category-level specialization, while MDMs maintain flatter activation profiles with fewer neurons strongly associated with any single semantic category.
4.2.3 Connecting Component-level and Neuron-level Perspectives
Component-wise logit lens and neuron-level visualization serve complementary interpretability roles. Logit lens probes semantic alignment at the level of components, revealing which tokens individual attention heads or MLPs are linearly aligned with and how they directly influence model outputs. Neuron-level analysis, by contrast, characterizes how semantic information is internally implemented, independent of direct alignment with the output vocabulary.
Together, these analyses reveal a consistent mechanistic pattern that is further supported by quantitative differences in explanation structure. ARMs exhibit a smaller number of unique explanatory components (266) but substantially higher variance in explanations (), indicating that model behavior is dominated by a limited set of highly influential and specialized components. In contrast, MDMs rely on a larger pool of explanatory components (426) with markedly lower variance (), suggesting that explanatory responsibility is distributed more evenly across components.
This quantitative shift aligns with the qualitative patterns observed in both analyses. ARMs employ a component-centric strategy in which task resolution is dominated by a small number of sharply specialized attention heads, or MLPs, supported by task-specific early-layer neurons. MDMs, by contrast, replace these sharp semantic roles with distributed responsibility, activating broader sets of components whose individual contributions are weaker but collectively stable. At the neuron level, this manifests as early-layer representations that exhibit reduced task-specific selectivity and more uniform activation profiles across tasks.
Taken together, these findings provide a coherent mechanistic account of diffusion post-training. MDMs trade component-level specialization for distributed semantic coverage, supported by early layers that provide broad, task-invariant representational scaffolding. This reorganization explains both the front-loaded computation observed in circuit analyses and the absence of dominant, easily interpretable components in logit-lens probes, highlighting a fundamental shift from localized causal pathways to globally integrated computation.
5 Conclusion
In this paper, we investigated how post-training an autoregressive language model (ARM) with a masked diffusion objective reshapes its internal computational mechanisms. By combining circuit analysis, component-wise logit-lens probing, and neuron-level explanaion, we provided a mechanistic comparison between ARMs and their post-trained masked diffusion model (MDM) counterparts across tasks with distinct structural demands.
Limitations
Our analysis focuses on a limited set of tasks—IOI and Countdown—which are chosen to isolate causal dependency and global planning behaviors, respectively. While these tasks are well-established benchmarks in mechanistic interpretability, the circuits identified in this work should be interpreted as task-specific mechanisms rather than universal properties of autoregressive or masked diffusion language models. Different linguistic or reasoning tasks may recruit alternative circuitry or induce distinct patterns of post-training adaptation. In addition, our circuit discovery pipeline relies on sparsity-inducing methods such as EAP-IG and Top-K divergence selection to isolate the most behaviorally salient components. Although this design choice improves interpretability by filtering out noisy or redundant pathways, it implies that not all attention heads and MLP components are exhaustively analyzed, and secondary or auxiliary mechanisms that contribute weakly to task performance may be omitted. Finally, our approach involves an inherent trade-off between interpretability resolution and computational scalability: component-wise logit lens analysis and neuron-level visualization across diffusion steps are computationally expensive, particularly for large-scale models. As a result, we prioritize in-depth analysis of selected components over exhaustive full-model sweeps, and extending the analysis to broader component sets or additional configurations would require substantially greater computational resources.
Appendix A Dataset Details
Datasets
For the countdown task, we evaluate both LLaMA-series and Qwen-series models using 500 evaluation examples per model. For LLaMA-series models, we use 13 diffusion steps, which correspond to 13 tokens (diffusion step = token). For Qwen-series models, we use 12 diffusion steps, corresponding to 12 tokens. For the IOI task, we evaluate 500 examples in a single-step setting with 1 diffusion step and 1 token.
Appendix B Experimental Details and Results
Computational Resources
All experiments were conducted on NVIDIA A6000 GPUs. Circuit extraction and analysis required less than 10 GPU-hours per model. No additional pre-training was performed.
Implementation Details
Our experiments use HuggingFace Transformers for model loading and inference. Circuit discovery is implemented using Edge Attribution Patching with Integrated Gradients (EAP-IG), where we select the top 1,000 edges(corresponding to a faithfulness score of 0.6) ranked by attribution scores as a trade-off between sparsity and attribution coverage. Empirically, smaller thresholds lead to unstable circuit structures, while larger thresholds rapidly approach the full graph without improving interpretability. The choice of 1,000 edges lies in a stable regime where circuit topology and component rankings remain consistent.
Attribution scores are then aggregated at the component level to identify the Top-K components. We set , corresponding to the smallest value for which the selected components form a well-connected circuit, ensuring that the resulting subgraph remains interpretable while preserving end-to-end connectivity. Logit lens projections follow the standard unembedding-based formulation. All other parameters follow default settings from the respective libraries.
Licenses and Terms of Use
All pretrained models and tools used in this work are publicly released research artifacts. We use them solely for research and analysis purposes, in accordance with their respective licenses, and do not redistribute any models or derived data.
Step-wise Circuit Stability
For analyses involving step-wise circuit extraction, we compute circuits at each diffusion step and then aggregate attribution scores across steps. We find that the set of participating components (attention heads and MLP layers) remains largely stable over diffusion time, with more than 90% overlap in selected components between steps. Accordingly, the step-wise circuit visualizations shown in Figures 6 and 5 represent averaged structures, while step-wise variation is primarily reflected in changes to the attributed edges rather than in the identity of the components themselves. This suggests that masked diffusion primarily refines information routing among a largely fixed component set, rather than progressively recruiting new components.
| Model | Top Source Components (Edges) |
| LLaMA-2 | inputm0, inputm1, a3.h26m3, a26.h21logits, m1a3.h26, a1.h18m1, inputa3.h26, a25.h0logits, m27logits, a27.h29logits, m0m1, m1a3.h26, a5.h15a6.h30, inputa3.h26, a24.h3logits, a23.h20logits, a21.h30logits, a1.h1m1, m2a3.h26, m0a3.h26, a18.h9logits, a21.h1logits, m2a3.h26, m4a6.h30, m0a3.h26, m0a1.h18, inputa1.h18, a1.h1a3.h26, a24.h15logits, a20.h8logits |
| Qwen | a24.h24logits, a26.h15logits, a23.h11logits, a27.h21logits, a27.h4logits, m27logits, a27.h1logits, a26.h26logits, a26.h22logits, m20m22, a27.h18logits, a27.h17logits, a27.h5logits, a27.h14logits, a24.h23logits, m24logits, inputa0.h10, a27.h3logits, a20.h24m22, a27.h24logits, a17.h24a20.h24, a26.h5logits, inputa0.h3, a27.h4m27, m25logits, a25.h24logits, a18.h25a20.h24, a18.h27a20.h24, a26.h2logits, m22a25.h25 |
| DiffuLLaMA | inputm0, inputm1, a3.h26m3, a28.h7logits, m31logits, a26.h21logits, m0m1, a1.h18m1, a27.h29logits, m1a3.h26, a1.h1m1, a5.h15a6.h30, m1a3.h26, m1m4, inputa3.h26, inputa3.h26, a26.h14logits, m2a3.h26, a30.h12logits, m4a6.h30, m0a3.h26, a23.h20logits, a18.h9a28.h7, inputa1.h18, m0a1.h18, inputm4, m2a3.h26, a22.h19logits, inputm2, a27.h29a28.h7 |
| Dream | a24.h24logits, m19m20, a23.h10logits, m18m19, m19m21, m9m19, m21m22, a18.h25m20, m14m15, m25logits, m12m18, m7m17, a15.h20m16, m9m20, m10m20, a15.h20m20, m12m15, m12m13, m18m22, m17m20, m14m22, a15.h20m19, a15.h23m16, m8m17, m11m16, m21logits, a25.h24logits, a18.h25m21, m8m20, m11m20 |
| Model | Top Source Components (Edges) |
| LLaMA-2 | inputa0.h15, m10m11, m11m12, a1.h22m1, m6m11, m11m15, inputa0.h25, m0a1.h22, m29logits, m7m9, m0m2, m28logits, m8m12, m8a11.h29, m8m11, a2.h2m3, m3m5, m14m15, m0m1, inputa0.h3, m13m15, a12.h5m13, m7m10, a12.h22m12, a5.h15a7.h6, inputa2.h2, m24logits, m0a1.h22, m12m14, m7a8.h15 |
| Qwen | m26logits, m25logits, m27logits, inputa0.h3, a23.h11logits, a25.h12logits, a26.h22logits, m24logits, a26.h24logits, m26m27, a24.h23logits, m21a23.h11, a0.h3m0, a22.h13logits, a26.h23logits, a23.h11m25, m25m27, a23.h19logits, a25.h12m27, a26.h25logits, m21a25.h12, m20a23.h11, a26.h26logits, m25a26.h22, a23.h11m26, a26.h22m27, a26.h22m26, a26.h24m27, a23.h11a26.h22, a24.h23m25 |
| DiffuLLaMA | m1m2, inputm0, m1m3, inputa0.h12, m31logits, m1a4.h5, inputa0.h3, inputa0.h15, m1a3.h3, m1a3.h27, inputm1, inputa0.h0, inputa0.h3, inputa0.h1, inputa1.h1, m1a3.h26, m1a3.h7, m1a3.h8, m1m4, m0m1, m1a4.h5, inputa0.h13, m1a2.h2, m0m3, m1a3.h0, inputa0.h3, m1a6.h20, inputa0.h13, m1a3.h17, m1a5.h23 |
| Dream | m25logits, m27logits, m26logits, inputa0.h15, a0.h10m0, inputa0.h3, m24logits, inputa0.h10, m23logits, inputa0.h11, inputa0.h15, inputa0.h15, a0.h3m0, a25.h12logits, m0m1, m26m27, inputa0.h0, inputa0.h11, a27.h11logits, a26.h22logits, a26.h25logits, m22logits, m21m27, a0.h15m0, inputa0.h11, m25m26, m22m27, a25.h12m27, inputa0.h10, a26.h24logits |
| Task | Model | Comp. | Top Tokens (Probability) |
| Countdown | DiffuLLaMA | m1 | sierp (0.936), kwiet (0.045), Hinweis (0.004) |
| m2 | (U+207B) (5.6e-05), nahm (5.6e-05), Hinweis (5.5e-05) | ||
| m3 | iftung (4.1e-05), (U+043D) (U+0434) (U+0434) (4.1e-05), iation (4.1e-05) | ||
| input | rd (3.7e-05), iation (3.7e-05), ness (3.7e-05) | ||
| m0 | mes (3.5e-05), led (3.5e-05), med (3.5e-05) | ||
| Dream | m25 | out (0.012), )) (0.001), from (0.001) | |
| m26 | ,- (7.9e-04), A (7.3e-04), S (6.3e-04) | ||
| m22 | nothing (8.7e-05), H (8.6e-05), a (7.2e-05) | ||
| m24 | int (8.3e-05), i (7.3e-05), commemor (6.5e-05) | ||
| m21 | = (3.2e-05), by (3.1e-05), C (3.0e-05) | ||
| a27.h11 | doen (2.9e-05), uteur (2.6e-05), retali (2.4e-05) | ||
| a25.h12 | 7 (2.2e-05), 8 (1.9e-05), 3 (1.7e-05) | ||
| a26.h22 | cosy (2.0e-05), W (1.9e-05), -ok (1.8e-05) | ||
| a26.h25 | fourth (1.4e-05), five (1.4e-05), IV (1.4e-05) | ||
| a26.h24 | ist (1.2e-05), /S (1.2e-05), SIX (1.2e-05) | ||
| m1 | . (1.0e-05), 1 (1.0e-05), # (1.0e-05) | ||
| m0 | e (9.0e-06), J (9.0e-06), G (9.0e-06) | ||
| a0.h10 | what (7.0e-06), that (7.0e-06), is (7.0e-06) | ||
| a0.h15 | one (7.0e-06), e (7.0e-06), the (7.0e-06) | ||
| a0.h3 | - (7.0e-06), in (7.0e-06), on (7.0e-06) | ||
| input | (7.0e-06), Increment (7.0e-06), 1 (7.0e-06) | ||
| LLaMA-2 | m1 | sierp (0.896), Unterscheidung (0.074), kwiet (0.027) | |
| m24 | them (0.009), ihnen (0.004), they (0.003) | ||
| m28 | — (0.006), , (0.005), – (0.004) | ||
| m14 | /- (7.7e-04), -+ (2.2e-04), ỳ (2.0e-04) | ||
| m15 | by (2.7e-04), look (2.1e-04), > (1.5e-04) | ||
| m13 | Halle (1.7e-04), Hook (1.2e-04), wa (1.1e-04) | ||
| m11 | attan (1.4e-04), iore (1.1e-04), (U+0442)(U+043A)(U+0443) (1.0e-04) | ||
| m12 | ieder (1.1e-04), ikai (1.1e-04), sail (1.1e-04) | ||
| m7 | bek (9.7e-05), untime (7.7e-05), mina (7.5e-05) | ||
| m9 | opsis (8.7e-05), PO (8.2e-05), zug (7.7e-05) | ||
| m10 | ador (8.4e-05), rok (8.3e-05), keit (8.3e-05) | ||
| m8 | concrete (8.3e-05), OST (8.1e-05), Chor (8.0e-05) | ||
| m6 | idenote (7.2e-05), ischof (7.1e-05), asm (7.1e-05) | ||
| m0 | bolds (6.6e-05), sce (6.0e-05), hina (5.7e-05) | ||
| m5 | kop (6.6e-05), Ő (6.3e-05), Sug (6.2e-05) | ||
| m2 | nobody (6.5e-05), nahm (6.0e-05), everybody (6.0e-05) | ||
| m3 | uche (5.2e-05), Chronology (5.1e-05), emer (4.9e-05) | ||
| a12.h5 | extension (3.9e-05), oba (3.9e-05), extensions (3.9e-05) | ||
| a5.h15 | Campbell (3.8e-05), beskre (3.7e-05), :(U+2009) (3.7e-05) | ||
| input | /- (3.7e-05), igny (3.6e-05), Extern (3.6e-05) | ||
| a1.h22 | (U+045A)y (3.4e-05), (U+4E0B) (3.4e-05), unci (3.4e-05) | ||
| a12.h22 | tel (3.4e-05), (3.3e-05), guez (3.3e-05) | ||
| a2.h2 | zik (3.2e-05), Muse (3.2e-05), èn (3.2e-05) | ||
| Qwen | m26 | (U+6027)(U+4EF7) (1.000), B (1.000), & (0.999) | |
| m27 | Human (1.000), ˆK (1.000), derive (0.999) | ||
| m24 | (U+62EC) (0.973), (U+5973)(U+6027)(U+670B)(U+53CB) (0.955), .ImageAlign (0.914) | ||
| m21 | aeda (0.608), so (0.599), to (0.306) | ||
| a26.h24 | A (0.032), A (0.006), (G (0.004) | ||
| a26.h23 | make (0.008), (make (0.008), .make (0.006) | ||
| a26.h26 | -await (4.0e-04), XMLElement (3.2e-04), etail (2.7e-04) | ||
[…]
| a26.h25 | (U+81EA)(U+52A8)(U+751F)(U+6210) (3.1e-04), /Dk (2.5e-04), line (2.4e-04) | ||
| a26.h22 | (U+5341)(U+56DB) (1.9e-04), (U+5341)(U+4E09) (1.8e-04), (U+80B2)(U+4EBA) (1.6e-04) | ||
| a25.h12 | 4 (1.5e-04), 5 (1.4e-04), Fifth (1.2e-04) | ||
| a23.h19 | Five (7.1e-05), five (5.0e-05), 5 (4.1e-05) | ||
| a24.h23 | num (2.5e-05), Gall (2.5e-05), num (2.4e-05) | ||
| a0.h3 | teenth (2.4e-05), bénéficie (1.9e-05), Noticed (1.9e-05) | ||
[…]
| input | (U+304D)(U+3061)(U+3093) (9.0e-06), (U+4E26)(U+4E14) (9.0e-06), (U+6362)(U+53E5)(U+8BDD) (9.0e-06) | ||
| IOI | DiffuLLaMA | m1 | sierp (0.550), kwiet (0.155), Hinweis (0.032) |
[…]
| a23.h20 | Sarah (3.8e-04), Vir (1.1e-04), sar (1.1e-04) | ||
| a28.h7 | V (2.9e-04), K (2.0e-04), Kim (2.0e-04) | ||
| a26.h14 | David (1.7e-04), David (1.4e-04), dav (1.2e-04) | ||
| a30.h12 | William (1.3e-04), Will (1.2e-04), Will (1.1e-04) | ||
| m4 | disambiguation (1.1e-04), printStackTrace (9.8e-05), - (9.5e-05) | ||
| m2 | nahm (5.5e-05), - (5.5e-05), Hinweis (5.4e-05) | ||
| a27.h29 | urn (5.4e-05), urr (5.3e-05), enis (4.9e-05) | ||
| a18.h9 | owo (5.1e-05), ceu (4.3e-05), fen (4.3e-05) | ||
| m3 | (U+4F1D) (3.5e-05), tu (3.5e-05), adr (3.5e-05) | ||
| a5.h15 | temps (3.4e-05), vend (3.4e-05), cancel (3.4e-05) | ||
| a3.h26 | lex (3.3e-05), ongodb (3.3e-05), huvudstaden (3.3e-05) | ||
| input | ô (3.3e-05), Horn (3.3e-05), roid (3.3e-05) | ||
| m0 | Einzeln (3.3e-05), (U+800C) (3.3e-05), atri (3.3e-05) | ||
| a1.h1 | (U+4EA4) (3.2e-05), gate (3.2e-05), Moc (3.2e-05) | ||
| a1.h18 | (3.2e-05), jsp (3.2e-05), epen (3.2e-05) | ||
| Dream | m21 | pliers (1.4e-05), Tap (1.4e-05), ynom (1.3e-05) | |
| m13 | doen (1.1e-05), bourgeois (1.1e-05), upholstery (1.0e-05) | ||
| m8 | wooded (1.1e-05), curt (1.0e-05), Genius (1.0e-05) | ||
| m9 | melodies (1.1e-05), interpolate (1.0e-05), Infantry (1.0e-05) | ||
| m10 | forestry (1.0e-05), rhet (1.0e-05), doen (1.0e-05) | ||
| m15 | orestation (1.0e-05), secluded (9.0e-06), cosy (9.0e-06) | ||
| m18 | Races (1.0e-05), weets (9.0e-06), ife (9.0e-06) | ||
| m19 | adjud (1.0e-05), enchanted (1.0e-05), instantiate (1.0e-05) | ||
| m20 | blot (1.0e-05), oval (1.0e-05), blinking (1.0e-05) | ||
| m7 | glimps (1.0e-05), seeding (1.0e-05), sadd (1.0e-05) | ||
| m11 | Tweet (9.0e-06), Intr (9.0e-06), enchanted (8.0e-06) | ||
| m12 | milit (9.0e-06), bourgeois (9.0e-06), slashes (9.0e-06) | ||
| m14 | enam (9.0e-06), upholstery (9.0e-06), vener (9.0e-06) | ||
| m16 | lan (9.0e-06), part (9.0e-06), ocal (9.0e-06) | ||
| m17 | Israelis (9.0e-06), commemor (9.0e-06), driv (9.0e-06) | ||
| a23.h10 | (8.0e-06), - (8.0e-06), 0 (8.0e-06) | ||
| a24.h24 | i (8.0e-06), in (8.0e-06), on (8.0e-06) | ||
| a15.h20 | ’s (7.0e-06), home (7.0e-06), half (7.0e-06) | ||
| a15.h23 | doen (7.0e-06), Packages (7.0e-06), classy (7.0e-06) | ||
| a18.h25 | ropy (7.0e-06), liner (7.0e-06), Bio (7.0e-06) | ||
| a25.h24 | mailed (7.0e-06), RSS (7.0e-06), masturbating (7.0e-06) | ||
| LLaMA-2 | m1 | sierp (0.865), Unterscheidung (0.110), kwiet (0.022) | |
| m27 | too (0.394), her (0.042), e (0.031) | ||
| a26.h21 | Marian (0.076), Pat (0.008), Anne (0.008) | ||
| a25.h0 | Richard (0.050), William (0.035), David (0.033) | ||
| a24.h3 | Rosa (0.007), Williams (6.4e-04), Alice (6.1e-04) | ||
| a20.h8 | Susan (0.004), sus (4.6e-04), suspect (1.4e-04) | ||
| a23.h20 | Sarah (0.004), Vir (0.001), vir (0.001) | ||
| a21.h30 | Lee (6.1e-04), Kelly (3.3e-04), ee (1.5e-04) | ||
| m4 | vy (2.1e-04), disambiguation (2.1e-04), - (2.0e-04) | ||
| a27.h29 | arta (1.3e-04), ML (1.3e-04), ignon (1.3e-04) | ||
| a18.h9 | Blue (9.5e-05), cyk (9.2e-05), nja (8.8e-05) | ||
| m0 | bolds (6.4e-05), sce (6.0e-05), partiellement (5.6e-05) | ||
| m2 | nobody (6.4e-05), nahm (5.9e-05), everybody (5.9e-05) | ||
| m3 | ime (5.4e-05), (U+82B1) (5.2e-05), ña (5.1e-05) | ||
| input | ny (4.0e-05), ten (4.0e-05), eral (3.9e-05) | ||
| a5.h15 | Chronology (3.7e-05), :// (3.6e-05), Extern (3.6e-05) | ||
| a3.h26 | erea (3.6e-05), ząt (3.6e-05), Songs (3.6e-05) | ||
| a1.h1 | (U+4EA4) (3.4e-05), Indep (3.3e-05), gate (3.3e-05) | ||
| a1.h18 | Bek (3.3e-05), arguments (3.3e-05), Millionen (3.3e-05) | ||
| Qwen | m27 | Human (1.000), Rossi (0.990), ‘‘ (0.978) | |
| m25 | Alexander (1.000), shall (0.986), zá (0.970) | ||
[…]
| m22 | thought (0.958), during (0.921), term (0.914) | ||
| m20 | (U+6027)(U+4EF7) (0.893), ",__ (0.247), ynos (0.101) | ||
| a27.h17 | Christina (0.444), Jessica (0.339), Crystal (0.330) | ||
| a27.h18 | Lisa (0.418), Elizabeth (0.282), Nic (0.228) | ||
| a27.h1 | Jamie (0.400), Nathan (0.345), Mary (0.275) | ||
| a27.h21 | Amy (0.360), Amber (0.331), Adam (0.331) | ||
| a26.h5 | Jesse (0.344), Nich (0.217), Rebecca (0.203) | ||
| a27.h3 | Katie (0.305), Ken (0.295), Brittany (0.189) | ||
| a27.h14 | Heather (0.252), Steven (0.244), Sean (0.223) | ||
| a27.h24 | Scott (0.233), Brad (0.221), Kris (0.220) | ||
| a26.h2 | Brad (0.227), Megan (0.194), brand (0.180) | ||
| a27.h4 | Mary (0.224), Ben (0.223), Mark (0.212) | ||
| a26.h15 | Danielle (0.215), Alicia (0.182), Dustin (0.176) | ||
| a24.h23 | John (0.013), Thomas (0.013), Kenneth (0.008) | ||
| a25.h24 | ch (0.002), William (0.001), w (0.001) | ||
| a24.h24 | Gad (0.001), ogen (9.4e-04), Aqu (8.7e-04) | ||
| a26.h22 | .AppSettings (6.5e-04), azt (6.0e-04), .d (4.3e-04) | ||
| a26.h26 | (U+0625)(U+0639)(U+062F)(U+0627)(U+062F) (1.9e-04), .TRAILING (1.8e-04), inx (1.8e-04) | ||
| a18.h25 | _locator (8.1e-05), ="’. (8.1e-05), .instrument (7.8e-05) | ||
| a20.h24 | setChecked (7.5e-05), anmar (6.5e-05), CAF (6.4e-05) | ||
| a18.h27 | (U+0623)(U+063A)(U+0644)(U+0628) (4.6e-05), dealloc (4.5e-05), (U+FFFD)(U+FFFD) (4.4e-05) | ||
| a17.h24 | .setCharacter (3.0e-05), -urlencoded (2.8e-05), entious (2.7e-05) | ||
| input | (U+304D)(U+3061)(U+3093) (9.0e-06), (U+6362)(U+53E5)(U+8BDD) (9.0e-06), (U+4E26)(U+4E14) (9.0e-06) |

### v3

@@
 Abstract
-Post-training pretrained Autoregressive models (ARMs) into Masked Diffusion models (MDMs) has emerged as a cost-effective strategy to overcome the limitations of sequential generation.
-However, the internal algorithmic transformations induced by this paradigm shift remain unexplored, leaving it unclear whether post-trained MDMs acquire genuine bidirectional reasoning capabilities or merely repackage autoregressive heuristics.
-In this work, we address this question by conducting a comparative circuit analysis of ARMs and their MDM counterparts.
-Our analysis reveals a systematic “mechanism shift” dependent on the structural nature of the task.
-Structurally, we observe a distinct divergence: while MDMs largely retain autoregressive circuitry for tasks dominated by local causal dependencies, they abandon initialized pathways for global planning tasks, exhibiting distinct rewiring characterized by increased early-layer processing.
-Semantically, we identify a transition from sharp, localized specialization in ARMs to distributed integration in MDMs.
-Through these findings, we conclude that diffusion post-training does not merely adapt model parameters but fundamentally reorganizes internal computation to support non-sequential global planning.
+Post-training pretrained autoregressive models (ARMs) into masked diffusion models (MDMs) has emerged as a cost-effective way to overcome the limitations of sequential generation.
+Yet the internal algorithmic changes induced by this shift remain poorly understood, leaving it unclear whether post-trained MDMs acquire genuine bidirectional reasoning or merely repackage autoregressive heuristics.
+We address this question through a comparative circuit analysis of ARMs and their MDM counterparts.
+Our analysis reveals a systematic “mechanism shift” that depends on the structural nature of the task.
+MDMs largely preserve autoregressive circuitry for tasks driven by local causal dependencies, but for global planning tasks they abandon initialized pathways and exhibit distinct rewiring with increased early-layer processing.
+At the semantic level, we observe a transition from sharp, localized specialization in ARMs to distributed integration in MDMs.
+These findings show that diffusion post-training does not simply adjust model parameters, but reorganizes internal computation to support non-sequential global planning.
 Mechanism Shift During Post-training from Autoregressive to Masked Diffusion Language Models
@@
 (2023).
-Despite these advances, the prevalent autoregressive framework Vaswani et al.
-(2017); Radford et al.
+Despite these advances, the prevalent autoregressive framework Radford et al.
 (2019) imposes structural limitations on generation Welleck et al.
@@
 (2025); Ye et al.
-(2025b).
+(2025c).
 Models such as Dream Ye et al.
-(2025b) demonstrate that this strategy can achieve strong performance while requiring only a fraction of the compute needed for training from scratch.
+(2025c) demonstrate that this strategy can achieve strong performance while requiring only a fraction of the compute needed for training from scratch.
 Despite the empirical success of post-training ARMs with diffusion objectives, the specific algorithm changes induced by this post-training process are not yet understood Gong et al.
@@
 (2025) to Dream Ye et al.
-(2025b), this line of work shows that post-training pretrained ARMs to MDM objectives can retain many practical advantages of diffusion—such as parallel decoding, iterative refinement, and bidirectional attention—while substantially reducing training cost.
+(2025c), this line of work shows that post-training pretrained ARMs to MDM objectives can retain many practical advantages of diffusion—such as parallel decoding, iterative refinement, and bidirectional attention—while substantially reducing training cost.
 Post-trained MDMs also achieve strong performance on directionality-sensitive tasks.
@@
 (2024).
-Across models and scales, similar circuits—such as induction heads—recur consistently, suggesting that they implement stable algorithmic functions rather than incidental patterns Prakash et al.
+Across models and scales, similar circuits,such as induction heads recur consistently, suggesting that they implement stable algorithmic functions rather than incidental patterns Prakash et al.
 (2024); Tigges et al.
@@
 (2025), comparable studies for text diffusion models remain limited.
-As a result, it is unclear whether diffusion objectives induce distinct computational strategies in language models or primarily reorganize existing autoregressive circuitry.
+As a result, it is unclear whether diffusion objectives induce distinct computational strategies in language models or reorganize existing autoregressive circuitry.
 3 Method
 To investigate the mechanistic shift from ARMs to MDMs, we adopt a comparative framework.
-Our primary objective is to distinguish the impact of shifting the learning objective from causal modeling to masked diffusion on circuit topology.
-We control for architectural confounding by analyzing ARMs alongside their directly post-trained MDM counterparts.
+We analyze pretrained ARMs alongside their directly post-trained MDM counterparts, allowing us to isolate the effect of changing the learning objective from causal modeling to masked diffusion while controlling for architectural confounding.
 3.1 Models and Configuration
@@
 (2023) architectures.
-To enable a direct comparison, we pair each standard autoregressive checkpoint with an MDM post-trained from that specific ARM base:
+To enable a direct comparison, we pair each autoregressive checkpoint with its corresponding post-trained MDM derived from the same backbone.
+Concretely, we compare Qwen2.5-7B with Dream-Base-7B Ye et al.
+(2025c), which is post-trained on the Qwen backbone, and LLaMA-2-7B with DiffuLLaMA-7B Gong et al.
+(2025), which is initialized from LLaMA 2 weights.
 3.2 Tasks and Datasets
-We analyze circuit behavior on two distinct tasks to highlight different performance capabilities: one focused on autoregressive causal dependencies and another on global reasoning.
-Additional details for tasks and datasets are provided in Appendix A.
+We select two tasks that lie at opposite ends of the local–global reasoning spectrum: one dominated by autoregressive causal dependencies and one requiring global constraint integration.
+Let denote the target token at position , the prefix context, and the full sequence context excluding .
+Using Dream-based conditional entropy, we define a task as exhibiting more local reasoning when , and more global reasoning when .
+Intuitively, the former indicates that prefix context is largely sufficient, whereas the latter indicates that bidirectional full-sequence context substantially reduces uncertainty.
+Additional details are provided in Appendix A.
 Indirect Object Identification (IOI):
@@
 (2023).
-We select this task to serve as a representative baseline for causal reasoning; specifically, we aim to investigate whether the induction circuitry—which inherently relies on sequential, left-to-right context to copy tokens—persists, dissolves, or transforms when the training objective shifts from causal prediction to masked diffusion.
+We use this task as a representative benchmark for local reasoning, asking whether induction circuitry—which relies on sequential, left-to-right context to copy tokens—is preserved, dissolved, or transformed when the learning objective shifts from causal prediction to masked diffusion.
 Countdown:
-A numerical reasoning task where the model is provided with a set of integers and a target value and must generate a valid arithmetic equation using the inputs to equal the target Ye et al.
+A numerical reasoning task in which the model is given a set of integers and a target value, and must generate a valid arithmetic expression that equals the target Ye et al.
 (2025a).
-This task serves as a critical benchmark for global reasoning because it requires inverse planning—decisions made early in the sequence must be strictly conditioned on the final goal—thereby challenging the left-to-right causality of ARMs while favoring the bidirectional context and global refinement capabilities of MDMs.
-Inference Configuration
-For all experiments, we use task-dependent generation lengths.
-For IOI, we use a single diffusion step, while for Countdown, we align diffusion steps with the target sequence length.
-Details are provided in Appendix A.
+We use this task as a representative benchmark for global reasoning, since it requires inverse planning: decisions made early in the sequence must be conditioned on the final goal.
+This structure challenges the left-to-right causality of ARMs while favoring the bidirectional context and iterative refinement capabilities of MDMs Ye et al.
+(2025b).
+Inference Configuration:
+We use task-dependent generation lengths in all experiments.
+For IOI, we use a single diffusion step; for Countdown, we align the number of diffusion steps with the target sequence length.
+Additional details are provided in Appendix A.
 3.3 Circuit Discovery and Analysis Pipeline
@@
 To verify this, we employ a three-stage pipeline: Discovery, Topological Comparison, and Mechanism Interpretation.
-This workflow allows us to first locate where the computation changes and then analyze what those changes represent semantically.
 3.3.1 Discovery
@@
 3.3.2 Attribution-Guided Circuit Comparison
-To localize where algorithmic changes concentrate when transitioning from ARMs to MDMs, we leverage EAP-IG to compare circuits at the level of edges and their incident components.
+To localize where algorithmic changes concentrate when transitioning from ARMs to MDMs, we leverage EAP-IG to compare circuits at the level of edges and the components they connect.
 Rather than treating all discovered edges equally, we focus on those that carry the highest attribution mass under EAP-IG and then identify which components are repeatedly used as sources or sinks of these high-attribution edges.
@@
 For each model (ARM and MDM), we first select the set of top-attribution edges (1000 edges), and , by ranking edges according to their EAP-IG scores on identical prompts.
-We then quantify overlap using the Jaccard similarity .
-High overlap suggests that the MDM reuses the same high-attribution pathways as the ARM, whereas low overlap indicates that diffusion training recruits a distinct set of edges to implement its generative computation.
+We quantify overlap using the Jaccard similarity .
+High overlap suggests that the MDM reuses the same pathways as the ARM, whereas low overlap indicates that diffusion training recruits a distinct set of edges to implement its generative computation.
 The choice of selecting the top 1000 edges reflects a trade-off between attribution coverage and circuit sparsity and is empirically justified in Appendix B.
@@
 To move from edges to components, we assign each node a score by aggregating the EAP-IG scores of all incident edges, defined as the sum of incoming attributions and outgoing attributions , where denotes the set of top-attribution edges for the model.
-In words, a component is important if it repeatedly appears as either the source or the target of high-scoring EAP-IG edges.
+Thus, a component is considered important if it repeatedly appears as either the source or the target of high-scoring EAP-IG edges.
 We then define the Top-K Components (with ) for each model as the nodes with the largest .
 The choice of is empirically motivated and discussed in Appendix B.
-Based on these metrics, we first assess the degree of circuit reuse between ARMs and MDMs via overlap of high-attribution edges.
+Layer-wise Center of Gravity (CoG):
+To quantify where the discovered computation is concentrated across depth, we compute a layer-wise Center of Gravity (CoG) from the EAP-IG attribution profile.
+Let denote the total EAP-IG attribution mass associated with layer in the discovered circuit, and define .
+Lower CoG indicates more front-loaded computation in earlier layers, while higher CoG indicates greater reliance on middle or late layers.
+We use CoG as a scalar summary of depth-wise circuit organization.
+Formal details are provided in Appendix C.
+Using these edge-, component-, and layer-level summaries, we first assess the degree of circuit reuse between ARMs and MDMs via overlap of high-attribution edges.
 High edge overlap provides evidence of mechanistic recycling, where the MDM relies on pre-trained autoregressive pathways, while low overlap indicates the emergence of diffusion-specific generative strategies.
-Within these regions, the Top-K EAP-IG Components pinpoint nodes that sit at the endpoints of the most influential edges, yielding a focused set of components for downstream mechanistic analysis via logit lens and neuron-level visualization.
+Among these edges, Top-K Components pinpoint the nodes at the most influential interactions, yielding a focused set for downstream analysis.
+CoG then provides a scalar summary of where this computation is depth-wise concentrated, complementing the overlap and component comparisons.
 3.3.3 Mechanism Interpretation
-Having identified where the circuit changes, we investigate how the computation differs by applying interpretability techniques specifically to the Top-K Divergent Components:
+Having identified where the circuit changes, we investigate how the computation differs by applying interpretability techniques specifically to the Top-K EAP-IG Components:
 Logit Lens and Component-wise Analysis:
-Following the logit lens framework nostalgebraist (2020), we project intermediate activations into the vocabulary space via the unembedding matrix , yielding and , where indexes a component (e.g., an attention head), denotes the token position, and the diffusion timestep.
+Following the logit lens framework nostalgebraist (2020), we project intermediate activations into vocabulary space via the unembedding matrix .
+Since each MDM is obtained by post-training from the same pretrained backbone as its paired ARM, we use as a shared linear readout basis across both models: , , where indexes a component, the token position, and the diffusion timestep.
+We treat these not as the model’s native predictive distributions, but as diagnostic probes of which vocabulary directions each component is aligned with under a shared semantic frame.
 For autoregressive models (Qwen, LLaMA), we apply the component-wise logit lens to all token positions in the sequence on both IOI and Countdown tasks.
-This enables us to trace how the model’s internal representation becomes aligned with the target output across the sequence and to decompose the contributions of the residual stream, attention heads, and MLPs at each token position.
+This enables us to trace how the model’s representation becomes aligned with the target output across the sequence and to decompose the contributions of the residual stream, attention heads, and MLPs at each token position.
 For MDMs (Dream, DiffuLLaMA), we examine how these component-wise projections evolve across the diffusion time steps at a fixed set of token positions.
@@
 Importantly, when applying the unembedding matrix to intermediate component activations, we do not interpret the resulting logits as the model’s final predictions.
-Instead, they serve as a diagnostic probe that reveals which tokens a component is linearly aligned with at a particular stage of computation.
+Instead, they serve as a diagnostic probe that reveals which tokens a component is linearly aligned with at a stage of computation.
+To complement qualitative inspection with scalar summaries of semantic specialization, we compute quantitative probes from the component-wise logit distributions.
+Specifically, we measure (i) NameFrac@K, the fraction of person-name tokens among the top- aligned tokens, where person-name tokens are identified using a pretrained BERT-based named entity recognition (NER) model filtered for the PERSON label; (ii) LogitGap, the gap between the top-1 and top-2 logits, which captures selective amplification; (iii) the log-mean-exp dominance gap (LME), which compares the dominance of name versus non-name tokens within the top- set; and (iv) top- entropy, which quantifies whether alignment is sharply concentrated or broadly distributed.
+Together, these probes provide an objective characterization of component-level specialization and prediction dominance.
+Formal definitions and details are provided in Appendix C.
 Neuron Explanation:
@@
 Following the methodology of Bills et al.
-(2023), we then extract the input tokens that elicit the largest-magnitude activations for a given neuron and use these highly activating examples for qualitative inspection and automated explanation.
-We focus this fine-grained analysis on the Top-K Divergent Components identified in the previous step: specifically, for divergent MLP layers, we visualize the top activating neurons to understand the specific attributes they extract; for divergent Attention layers, we analyze the head’s output features to determine what information is being moved.
+(2023), we then extract the input tokens that elicit the largest-magnitude activations for a given neuron and use these activating examples for qualitative inspection and automated explanation.
+We focus this fine-grained analysis on the Top-K EAP-IG components identified in the previous step: specifically, for high-attribution MLP layers, we visualize the top activating neurons to understand the specific attributes they extract; for high-attribution Attention layers, we analyze the head’s output features to determine what information is transmitted.
 The primary goal of this visualization is to characterize the semantic feature selectivity of the circuit.
-By mapping high-activation neurons to their corresponding tokens, we aim to determine if MDM components have learned to encode non-causal features (e.g., attending to or encoding “future target” tokens available during the diffusion process) that are fundamentally inaccessible to the standard autoregressive model.
+By mapping high-activation neurons to their corresponding tokens, we aim to determine if MDM components have learned to encode non-causal features (e.g., attending to or encoding “future target” tokens available during the diffusion process) that are fundamentally inaccessible to the ARMs.
+Together with the component-level logit-lens probes, this neuron-level analysis helps distinguish whether diffusion post-training changes only the distribution of output-aligned predictions or also reshapes the underlying internal features that support those predictions.
 4 Results & Analysis
 4.1 Circuit-Level Differences
-| Setting | Edge Overlap | Top- Overlap |
-| IOI (Qwen / Dream) | 0.193 | 0.105 |
-| IOI (LLaMA-2 / DiffuLLaMA) | 0.088 | 0.124 |
-| Countdown (Qwen / Dream) | 0.008 | 0.093 |
-| Countdown (LLaMA-2 / DiffuLLaMA) | 0.032 | 0.081 |
-Figure 2 illustrates the circuit structures extracted for both ARMs and MDMs across the IOI and Countdown tasks, where each circuit represents the average structure aggregated over multiple prompts and diffusion steps.
-For the IOI task, the overall circuit topology of Dream closely resembles that of Qwen, with dominant interactions concentrated in similar layers and mediated by comparable attention pathways.
-A similar pattern is observed for DiffuLLaMA relative to its LLaMA-2 counterpart.
-This indicates that, for tasks primarily driven by local or causal dependencies—such as IOI, which effectively requires only shallow or single-step generation—post-trained MDMs largely preserve the autoregressive circuitry inherited from their ARM initializations.
-This qualitative similarity is further supported by quantitative circuit similarity metrics reported in Table 1.
-Across both model families, the IOI task exhibits consistently higher overlap in high-attribution edges and Top- components compared to the Countdown task, indicating substantial mechanistic reuse rather than the emergence of new diffusion-specific circuitry in tasks dominated by local dependency structures.
-In contrast, the Countdown task exhibits a clear shift in circuit organization between ARMs and MDMs.
-In the Countdown task, the ARM circuit becomes increasingly dense toward mid-to-late layers, whereas the MDM circuit concentrates a larger fraction of interactions in early layers, resulting in a front-loaded organization of computation.
-This redistribution is consistently observed both in the aggregated circuit statistics (Figure 2) and in the representative circuit visualizations (Figure 3), indicating a systematic reallocation of computation under the diffusion objective.
-This depth-wise shift toward increased early-layer engagement is consistently observed across both model families, with Dream and DiffuLLaMA exhibiting increased early-layer engagement compared to their respective autoregressive baselines.
-Quantitatively, this divergence is reflected in markedly lower edge-level and component-level overlap scores for the Countdown task (Table 1), indicating that diffusion post-training induces non-trivial reorganization of the underlying circuitry.
-Together, these results reveal a clear dichotomy in how masked diffusion post-training reshapes internal circuitry.
-For tasks dominated by local, causal dependencies such as IOI, post-trained MDMs retain a substantial fraction of the original autoregressive circuitry.
-High overlap in both edge-level attribution and Top- components indicates that the causal reasoning machinery learned during autoregressive pre-training remains functionally effective and is therefore reused rather than replaced.
-In contrast, for tasks requiring global constraint satisfaction, such as Countdown, the autoregressive circuitry proves insufficient.
-Here, diffusion post-training induces the emergence of distinct early-layer components that are largely absent from the original ARM circuits, accompanied by a marked collapse in circuit overlap.
-These newly emphasized early-layer components become the primary substrates for computation, suggesting that MDMs restructure their internal mechanisms to support global planning by front-loading computation rather than relying on sequential, causal pathways.
-Taken together, these findings demonstrate that masked diffusion post-training does not simply overwrite autoregressive mechanisms.
-Instead, it selectively preserves causal circuitry where it remains compatible with the task, while inducing genuinely new, diffusion-specific components when global reasoning demands exceed the expressive capacity of autoregressive computation.
+Figure 2 shows the circuits for ARMs and MDMs on IOI and Countdown. Combined with the overlap and CoG metrics in Table 1, the results reveal a task-dependent dichotomy.
+For IOI, MDMs largely preserve the autoregressive circuitry inherited from their ARM initializations.
+Table 1 quantifies this structural retention using both edge overlap and CoG.
+Although the absolute overlap score appears numerically small (e.g., 0.193 for Qwen/Dream), causal ablation shows that the shared edges form the functional core of the circuit: ablating the intersection reduces IOI accuracy to 28.1%, whereas ablating a random subset of non-shared edges of equal size retains 92.4% ( 1.2%) accuracy across five random seeds.
+The CoG also remains relatively stable in the mid-to-late layers (17.5 vs.
+20.4), indicating that causal reasoning machinery learned during pre-training is reused rather than replaced.
+| Task & Models | Overlap | CoG (Layer) | ||
+| Edge | Top- | ARM | MDM | |
+| IOI Qwen / Dream | 0.193 | 0.105 | 17.5 | 20.4 |
+| IOI LLaMA / DiffuLLaMA | 0.088 | 0.124 | 18.2 | 19.8 |
+| Countdown Qwen / Dream | 0.008 | 0.093 | 16.5 | 4.8 |
+| Countdown LLaMA / DiffuLLaMA | 0.032 | 0.081 | 17.1 | 5.3 |
+In contrast, for Countdown, the autoregressive circuitry is no longer sufficient.
+Circuit overlap drops sharply (e.g., to 0.008 for Qwen/Dream), indicating substantial structural decoupling.
+This change is accompanied by a strong shift in depth, with the CoG moving from layer 16.5 in the ARM to 4.8 in the MDM.
+This depth shift is also consistent with the aggregated layer-wise component usage statistics, which show increased concentration in earlier layers (Appendix B Figure 3).
+Taken together, these findings show that masked diffusion post-training does not simply overwrite autoregressive mechanisms.
+Instead, it preserves causal circuitry where it remains task-compatible, while inducing new early-layer diffusion components when global reasoning exceeds the limits of sequential autoregressive computation.
 4.2 Semantic Reorganization of Task-Critical Components and Early Layers
+| Category | Dream | DiffuLLaMA | Qwen | LLaMA |
+| PERSON | 1,300 | 337 | 298 | 528 |
+| OTHER | 8,348 | 3,251 | 2,965 | 14,606 |
+| ORG | 62 | 30 | 81 | 418 |
+| LOC | 55 | 33 | 31 | 343 |
+| DIGIT | 32 | – | 2 | 55 |
+| OPERATOR | 20 | 2 | 210 | 124 |
+| MISC | 3 | 2 | 2 | – |
+| Total | 9,820 | 3,933 | 3,775 | 15,619 |
+| NameFrac@10 | 13.2% | 8.6% | 7.9% | 3.4% |
+To understand how circuit-level differences translate into concrete computational strategies, we analyze model behavior at two complementary granularities.
+First, we examine task-critical components—attention heads and MLPs that receive high EAP-IG attribution—using component-wise logit lens analysis.
+This allows us to characterize how semantic roles are assigned to components that directly influence task outputs.
+Second, we analyze early-layer neuron activations to understand how these semantic roles are implemented internally in regions where MDMs emphasize computation but component-level semantics appear diffuse.
+4.2.1 Task-Critical Components: Component-wise Logit Lens Analysis
 | Task | Model | Component | Mean Logit | Top Tokens | Role |
@@
 | DiffuLLaMA | m4 | 1.40 | – | Symbol-related Component |
-To understand how circuit-level differences translate into concrete computational strategies, we analyze model behavior at two complementary granularities.
-First, we examine task-critical components—attention heads and MLPs that receive high EAP-IG attribution—using component-wise logit lens analysis.
-This allows us to characterize how semantic roles are assigned to components that directly influence task outputs.
-Second, we analyze early-layer neuron activations to understand how these semantic roles are implemented internally, particularly in regions where MDMs emphasize computation but component-level semantics appear diffuse.
-4.2.1 Task-Critical Components: Component-wise Logit Lens Analysis
-We begin by analyzing components that directly contribute to task performance, as identified by high EAP-IG attribution.
-By applying a component-wise logit lens to these components, we probe their semantic alignment with output tokens, allowing us to assess whether task-relevant information is concentrated within a small number of specialized components or distributed across many.
-Causal reasoning task (IOI)
-For the IOI task, ARMs exhibit sharply localized semantic specialization.
-In both Qwen and LLaMA, a small number of attention heads are strongly aligned with person-name tokens (e.g., Dan, Jerry), producing large-magnitude logits that dominate prediction.
-This pattern indicates that IOI is largely solved by a limited set of highly specialized components that act as deterministic, pointer-like mechanisms.
-DiffuLLaMA largely preserves this behavior.
-Several high-attribution components remain strongly aligned with person-name tokens and exhibit logit profiles comparable to those of the underlying ARM.
-This suggests substantial inheritance of autoregressive circuitry, consistent with the observation that IOI relies primarily on local causal dependencies that are already well supported by the autoregressive objective.
-Dream, however, departs from this pattern.
-Its high-attribution components do not exhibit clear person-name specificity.
-Instead, the strongest logit alignments correspond to non-person proper nouns or broader semantic tokens (e.g., Browser, Kremlin).
-This indicates that, despite similar circuit topology at a coarse level, diffusion post-training redistributes semantic roles across components, weakening the dominance of individual name-specific heads.
-Global reasoning task (Countdown).
-A contrasting pattern emerges in the Countdown task.
-ARMs rely on components with strong numerical selectivity, often sharply aligned with specific digits or operators.
-These components produce high-confidence logits concentrated on a narrow subset of tokens, indicating that global planning is approximated through sequential, component-centric heuristics.
-In Dream, this sharp specialization collapses.
-Instead of a single dominant numerical component, multiple components exhibit moderate logit responses distributed across numerical tokens, with no individual component exerting decisive control over prediction.
-DiffuLLaMA again occupies an intermediate regime, retaining numerical associations but with reduced magnitude and increased dispersion.
+We analyze task-critical components identified by high EAP-IG attribution using a component-wise logit lens.
+This allows us to test whether task-relevant information is concentrated in a few specialized components or distributed across many.
+We report probes of semantic alignment and dominance in the main text, and defer representative top-token examples to Table 3, LABEL:tab:logit_lens.
+Causal reasoning task (IOI):
+For IOI, ARMs exhibit sharply localized semantic specialization.
+This pattern appears in the way person-name tokens are concentrated among the aligned outputs of task-critical components.
+As shown in Table 2, MDMs can place person-name tokens relatively often among their top aligned candidates—Dream achieves a higher (13.2%) than Qwen (7.9%) and LLaMA (3.4%), while DiffuLLaMA reaches 8.6%—but this does not imply sharper specialization.
+Instead, ARMs exhibit substantially stronger selective amplification once a relevant name token appears.
+Qwen and LLaMA achieve median LME values of 0.90 and 0.89, respectively, compared with 0.31 for Dream and 0.05 for DiffuLLaMA (Appendix Table 9).
+ARMs also show larger logit gaps, while DiffuLLaMA exhibits higher entropy, consistent with a more distributed alignment pattern.
+In short, raw name frequency does not imply sharper specialization: ARMs remain dominated by a few decisive name-specific components, whereas MDMs spread name-related evidence across a broader set of components.
+This qualitative difference is also visible in representative component-level examples (Table 3).
+In Qwen and LLaMA, several high-attribution attention heads align strongly with person-name tokens, consistent with highly specialized, pointer-like behavior.
+DiffuLLaMA preserves this pattern, indicating substantial inheritance of autoregressive circuitry.
+Dream, however, departs from it: its high-attribution components are less clearly name-specific, and their strongest alignments often correspond to non-person proper nouns or broader semantic tokens.
+Thus, even when circuit topology is similar, diffusion post-training can redistribute semantic roles across components and weaken the dominance of individual name-specific heads.
+Global reasoning task (Countdown):
+A contrasting pattern emerges in Countdown.
+ARMs rely on components with strong numerical selectivity, sharply aligned with specific digits or operators, suggesting that global planning is approximated through sequential, component-centric heuristics (Table 3).
+In Dream, this sharp specialization collapses: instead of a single dominant numerical component, multiple components exhibit moderate responses across numerical tokens with no individual component exerting decisive control.
+DiffuLLaMA occupies an intermediate regime, retaining numerical associations but with reduced magnitude and increased dispersion.
+This redistribution is also architectural—Dream’s CoG on Countdown drops from 16.5 (Qwen) to 4.8 (Table 1), indicating that diffusion post-training front-loads task-relevant computation toward earlier layers.
 Taken together, these results indicate that diffusion post-training induces a semantic reorganization at the component level.
-Whereas ARMs resolve tasks through a small number of highly specialized components, MDMs—particularly on global reasoning tasks—redistribute semantic responsibility across multiple components, yielding a more ensemble-like computation.
+Whereas ARMs resolve tasks through a small number of highly specialized components, MDMs redistribute semantic responsibility across multiple components, yielding a more ensemble-like computation—accompanied in Dream by a pronounced shift toward earlier layers.
 4.2.2 Early-layer Representation: Neuron-level Analysis
-While component-wise logit lens analysis captures semantic specialization among task-critical components, it does not fully explain the behavior of early-layer regions that are emphasized by masked diffusion circuits.
-To characterize how task-relevant information is implemented in these layers, we analyze neuron activations in the lowest transformer layers across tasks and model classes.
-In ARMs, neuron activation in early layers is highly task-dependent.
-In the IOI task, a large number of neurons in early layers are strongly activated by descriptive adjectives and modifier-related tokens, suggesting that early layers encode fine-grained syntactic and descriptive features prior to resolving entity identity.
-In the Countdown task, this pattern shifts markedly toward neurons associated with numerical or technical content.
+While component-wise logit lens analysis captures semantic specialization among task-critical components, it does not fully explain the behavior of early-layer regions emphasized by MDM circuits.
+To characterize how task-relevant information is implemented there, we analyze neuron activations in the lowest transformer layers.
+This analysis is especially important for Countdown, where the CoG shift in Table 1 indicates that post-training moves substantial computation into the bottom of the network.
+In ARMs, early-layer neuron activation is highly task-dependent.
+In IOI, many early-layer neurons are strongly activated by descriptive adjectives and modifier-related tokens (e.g., 2,821 neurons in LLaMA; Appendix Table 5), suggesting these layers encode syntactic features before resolving entity identity.
+In Countdown, this pattern shifts toward neurons associated with numerical or technical content (over 1,000 neurons).
 Despite this task-dependent reallocation, ARMs consistently exhibit strong concentration within specific semantic categories.
 MDMs display a qualitatively different pattern.
-Across both IOI and Countdown, the total number of active neurons in early layers remains relatively stable.
-Moreover, the activated neurons tend to correspond to broad, genre-level cues—such as general numerical or technical content—rather than sharply defined task-specific categories.
-This indicates that MDMs rely less on early-layer specialization and instead maintain a more uniform, task-agnostic representational regime.
-This contrast persists in the early layer.
-Across depth, ARMs repeatedly exhibit sharper category-level specialization, while MDMs maintain flatter activation profiles with fewer neurons strongly associated with any single semantic category.
+Across both tasks, the number of active neurons in early layers remains relatively stable (typically 50–200 active neurons per top category; Appendix Table 5), and these neurons tend to correspond to broad, genre-level cues rather than sharply defined task-specific categories.
+This suggests MDMs rely less on category-specific early-layer specialization and instead maintain a more uniform, task-agnostic representational regime.
+Combined with the CoG results, this pattern implies that MDM early layers absorb more computation without relying on sharply specialized neurons.
 4.2.3 Connecting Component-level and Neuron-level Perspectives
-Component-wise logit lens and neuron-level visualization serve complementary interpretability roles.
-Logit lens probes semantic alignment at the level of components, revealing which tokens individual attention heads or MLPs are linearly aligned with and how they directly influence model outputs.
+Taken together, the component-level and neuron-level analyses reveal a consistent mechanistic pattern.
+Component-wise logit lens shows how task-relevant outputs are distributed across influential components, whereas neuron-level analysis shows how these representations are implemented in the early layers.
+More specifically, logit lens probes semantic alignment at the component level, revealing which tokens individual attention heads or MLPs are linearly aligned with and how they influence model outputs.
 Neuron-level analysis, by contrast, characterizes how semantic information is internally implemented, independent of direct alignment with the output vocabulary.
-Together, these analyses reveal a consistent mechanistic pattern that is further supported by quantitative differences in explanation structure.
-ARMs exhibit a smaller number of unique explanatory components (266) but substantially higher variance in explanations (), indicating that model behavior is dominated by a limited set of highly influential and specialized components.
-In contrast, MDMs rely on a larger pool of explanatory components (426) with markedly lower variance (), suggesting that explanatory responsibility is distributed more evenly across components.
-This quantitative shift aligns with the qualitative patterns observed in both analyses.
-ARMs employ a component-centric strategy in which task resolution is dominated by a small number of sharply specialized attention heads, or MLPs, supported by task-specific early-layer neurons.
+In ARMs, explanatory mass is concentrated in a few highly influential components, whereas in MDMs it is distributed across a broader pool of weaker but collectively supportive components.
+ARMs exhibit a smaller number of unique explanatory components (266) but a higher standard deviation in explanations (), indicating that model behavior is dominated by a limited set of highly influential and specialized components.
+In contrast, MDMs rely on a larger pool of explanatory components (426) with lower standard deviation (), suggesting that explanatory responsibility is distributed more evenly across components.
+ARMs employ a component-centric strategy in which task resolution is dominated by a small number of sharply specialized attention heads or MLPs, supported by task-specific early-layer neurons.
 MDMs, by contrast, replace these sharp semantic roles with distributed responsibility, activating broader sets of components whose individual contributions are weaker but collectively stable.
-At the neuron level, this manifests as early-layer representations that exhibit reduced task-specific selectivity and more uniform activation profiles across tasks.
-Taken together, these findings provide a coherent mechanistic account of diffusion post-training.
+At the neuron level, this appears as early-layer representations with reduced task-specific selectivity and more uniform activation profiles across tasks.
+Taken together, these findings provide a mechanistic account of diffusion post-training.
 MDMs trade component-level specialization for distributed semantic coverage, supported by early layers that provide broad, task-invariant representational scaffolding.
-This reorganization explains both the front-loaded computation observed in circuit analyses and the absence of dominant, easily interpretable components in logit-lens probes, highlighting a fundamental shift from localized causal pathways to globally integrated computation.
+This reorganization explains both the front-loaded computation observed in circuit analyses and the reduced dominance of interpretable components in logit-lens probes, highlighting a fundamental shift from localized causal pathways to more globally integrated computation.
 5 Conclusion
-In this paper, we investigated how post-training an autoregressive language model (ARM) with a masked diffusion objective reshapes its internal computational mechanisms.
-By combining circuit analysis, component-wise logit-lens probing, and neuron-level explanaion, we provided a mechanistic comparison between ARMs and their post-trained masked diffusion model (MDM) counterparts across tasks with distinct structural demands.
+In this paper, we investigated where and how masked diffusion post-training reshapes the internal mechanisms of ARMs.
+To identify where, we use circuit analysis and layer-wise CoG to trace changes in circuit composition and depth-wise computation.
+To understand how, we use component-wise logit-lens probing and neuron-level analysis to examine shifts in semantic role assignment and representation patterns.
+Our results show that masked diffusion post-training induces a task-dependent mechanism shift rather than merely reusing autoregressive computation.
+In IOI, MDMs largely preserve the core autoregressive circuit while replacing sharp specialization with distributed semantic responsibility.
+In Countdown, MDMs show broader circuit rewiring, a stronger shift toward earlier layers, and the same transition to distributed component roles.
+Overall, masked diffusion post-training preserves autoregressive mechanisms for local causal reasoning, while reorganizing them for more globally integrated, non-sequential reasoning.
 Limitations
-Our analysis focuses on a limited set of tasks—IOI and Countdown—which are chosen to isolate causal dependency and global planning behaviors, respectively.
-While these tasks are well-established benchmarks in mechanistic interpretability, the circuits identified in this work should be interpreted as task-specific mechanisms rather than universal properties of autoregressive or masked diffusion language models.
-Different linguistic or reasoning tasks may recruit alternative circuitry or induce distinct patterns of post-training adaptation.
-In addition, our circuit discovery pipeline relies on sparsity-inducing methods such as EAP-IG and Top-K divergence selection to isolate the most behaviorally salient components.
-Although this design choice improves interpretability by filtering out noisy or redundant pathways, it implies that not all attention heads and MLP components are exhaustively analyzed, and secondary or auxiliary mechanisms that contribute weakly to task performance may be omitted.
-Finally, our approach involves an inherent trade-off between interpretability resolution and computational scalability: component-wise logit lens analysis and neuron-level visualization across diffusion steps are computationally expensive, particularly for large-scale models.
-As a result, we prioritize in-depth analysis of selected components over exhaustive full-model sweeps, and extending the analysis to broader component sets or additional configurations would require substantially greater computational resources.
-Appendix A Dataset Details
+Our study is confined to two benchmark tasks, IOI and Countdown, selected to capture two core yet contrasting modes of computation central to our inquiry: locally causal dependency tracking and global sequence-level planning.
+This pairing enables a controlled comparison of whether post-trained masked diffusion models preserve autoregressive mechanisms or develop distinct computational circuitry under differing structural demands.
+Nevertheless, the circuits identified here should be understood as task-specific findings rather than universal signatures of autoregressive or masked diffusion language models.
+Establishing broader generality will require evaluation across additional tasks spanning diverse compositional, linguistic, and reasoning structures.
+Furthermore, while our attribution-guided pipeline enables targeted mechanistic analysis by isolating the most behaviorally salient edges and components, it does not constitute an exhaustive characterization of all computations underlying model behavior.
+A more complete account would necessitate wider circuit sweeps across a broader range of tasks, models, and training conditions — an endeavor that remains computationally prohibitive at scale.
+Appendix A Task Details
 Datasets
-For the countdown task, we evaluate both LLaMA-series and Qwen-series models using 500 evaluation examples per model.
-For LLaMA-series models, we use 13 diffusion steps, which correspond to 13 tokens (diffusion step = token).
-For Qwen-series models, we use 12 diffusion steps, corresponding to 12 tokens.
-For the IOI task, we evaluate 500 examples in a single-step setting with 1 diffusion step and 1 token.
-Appendix B Experimental Details and Results
+For Countdown, we evaluate both LLaMA-series and Qwen-series models on 500 examples per model.
+LLaMA-series models use 13 diffusion steps (13 tokens), while Qwen-series models use 12 diffusion steps (12 tokens).
+For IOI, we evaluate 500 examples in a single-step setting with 1 diffusion step and 1 token.
+Since IOI in our setup requires generating only a single target token, increasing the number of diffusion steps would still correspond to repeatedly refining the same one-token prediction.
+As a result, the underlying circuit is not expected to change qualitatively across multiple steps, making the single-step setting the most direct and interpretable choice.
+Dream-based conditional-entropy criterion
+Following Section 3, we operationalize the local–global reasoning criterion using Dream, which yields the most stable and interpretable conditional-entropy estimates among the MDMs we study.
+For a target token , we compare prefix-only entropy, , with full-context entropy excluding the target token, .
+A larger positive gap indicates stronger global reasoning, while a smaller or negative gap indicates a more local or prefix-dominant regime.
+Entropy-gap results
+Table 4 reports Dream-based conditional entropy for the two tasks.
+Countdown shows a positive entropy gap, indicating that full-sequence context substantially reduces uncertainty relative to prefix-only conditioning.
+In contrast, IOI shows a negative gap, consistent with a more local or prefix-dominant regime.
+| Task | Prefix-only | Full-context | Gap |
+| Countdown | 1.2756 [1.2517, 1.2984] | 0.8552 [0.7879, 0.9202] | +0.4204 [0.3456, 0.4911] |
+| IOI | 3.4383 [3.3125, 3.5644] | 4.1884 [4.0699, 4.3050] | -0.7500 [-0.9237, -0.5783] |
+Appendix B Experimental Details and Additional Results
 Computational Resources
@@
 Implementation Details
-Our experiments use HuggingFace Transformers for model loading and inference.
-Circuit discovery is implemented using Edge Attribution Patching with Integrated Gradients (EAP-IG), where we select the top 1,000 edges(corresponding to a faithfulness score of 0.6) ranked by attribution scores as a trade-off between sparsity and attribution coverage.
-Empirically, smaller thresholds lead to unstable circuit structures, while larger thresholds rapidly approach the full graph without improving interpretability.
-The choice of 1,000 edges lies in a stable regime where circuit topology and component rankings remain consistent.
+We use HuggingFace Transformers for model loading and inference.
+Circuit discovery is implemented with Edge Attribution Patching with Integrated Gradients (EAP-IG).
+We select the top 1,000 edges, corresponding to a faithfulness score of 0.6, as a trade-off between sparsity, attribution coverage, and graph connectivity.
+In practice, reducing the threshold below 1,000 leads to fragmented circuits: we observe that at 800 edges, the discovered subgraph no longer preserves end-to-end connectivity.
+By contrast, increasing the threshold yields graphs that more closely approach the full computation graph without providing additional interpretive benefit.
+The choice of 1,000 edges therefore lies in a stable regime where circuit topology, connectivity, and component rankings remain consistent.
 Attribution scores are then aggregated at the component level to identify the Top-K components.
-We set , corresponding to the smallest value for which the selected components form a well-connected circuit, ensuring that the resulting subgraph remains interpretable while preserving end-to-end connectivity.
-Logit lens projections follow the standard unembedding-based formulation.
-All other parameters follow default settings from the respective libraries.
+We set , the smallest value for which the selected components form a well-connected and interpretable subgraph while preserving end-to-end connectivity.
+Importantly, these top 100 components account for more than 65% of the total attribution mass, indicating that a relatively small subset of components captures most of the task-relevant computation.
+Logit-lens projections follow the standard unembedding-based formulation.
+All other parameters use default library settings.
 Licenses and Terms of Use
 All pretrained models and tools used in this work are publicly released research artifacts.
-We use them solely for research and analysis purposes, in accordance with their respective licenses, and do not redistribute any models or derived data.
+We use them solely for research and analysis in accordance with their respective licenses, and do not redistribute any models or derived data.
 Step-wise Circuit Stability
-For analyses involving step-wise circuit extraction, we compute circuits at each diffusion step and then aggregate attribution scores across steps.
-We find that the set of participating components (attention heads and MLP layers) remains largely stable over diffusion time, with more than 90% overlap in selected components between steps.
-Accordingly, the step-wise circuit visualizations shown in Figures 6 and 5 represent averaged structures, while step-wise variation is primarily reflected in changes to the attributed edges rather than in the identity of the components themselves.
-This suggests that masked diffusion primarily refines information routing among a largely fixed component set, rather than progressively recruiting new components.
+For step-wise circuit extraction, we compute circuits at each diffusion step and aggregate attribution scores across steps.
+The set of participating components (attention heads and MLP layers) remains largely stable over diffusion time, with more than 70% overlap in selected components between steps.
+Accordingly, the visualizations in Figures 7 and 6 represent averaged structures; step-wise variation is reflected mainly in the attributed edges rather than in component identity.
+This suggests that masked diffusion primarily refines information routing within a largely fixed component set, rather than progressively recruiting new components.
+| Task | Model | Layer | Explanation | Count |
+| IOI | LLaMA | 0–4 | descriptive adjectives | 2821 |
+| expressions of uncertainty or doubt | 104 | |||
+| synonyms for happiness or joy | 50 | |||
+| DiffuLLaMA | 0–4 | synonyms for happiness or joy | 193 | |
+| expressions of uncertainty or doubt | 80 | |||
+| technical jargon or specialized terminology | 57 | |||
+| Countdown | LLaMA | 0–4 | numerical values or quantities | 1012 |
+| terms associated with technology and innovation | 226 | |||
+| technical terms related to technology | 86 | |||
+| DiffuLLaMA | 0–4 | synonyms for happiness or joy | 207 | |
+| expressions of uncertainty or doubt | 114 | |||
+| terms associated with technology and innovation | 72 |
 | Model | Top Source Components (Edges) |
@@
 | input | (U+304D)(U+3061)(U+3093) (9.0e-06), (U+6362)(U+53E5)(U+8BDD) (9.0e-06), (U+4E26)(U+4E14) (9.0e-06) |
+Appendix C Quantitative Mechanism Metrics & Results
+To formalize specialization and dominance in our component-level logit lens analysis, and to provide objective support for our qualitative interpretations, we introduce several quantitative probes.
+Let denote the logit assigned by component to token , and let denote the set of top- tokens ranked by logit.
+C.1 Semantic Alignment and Dominance Probes
+Name Alignment Frequency (NameFrac@K):
+To quantitatively assess whether semantic alignment extends beyond the top-1 token, we measure the fraction of person-name tokens among the top- aligned tokens for each component:
+where denotes the set of person-name tokens, identified using a pretrained BERT-based Named Entity Recognition (NER) model (filtered for the PERSON label).
+Selective Amplification (Logit Gap):
+To quantify whether a component strongly favors a single token, we compute the logit gap between the highest and second-highest logits.
+Because softmax probabilities depend exponentially on logit differences, larger logit gaps indicate stronger selective amplification:
+where and represent the maximum and second-maximum logits, respectively.
+Log-Mean-Exp Dominance Gap (LME):
+Because frequency alone does not measure predictive dominance, we quantify whether name tokens exert stronger influence when they appear by computing a log-mean-exp dominance gap between name and non-name tokens within the top- set:
+where the LME for a subset of tokens is defined as:
+This metric controls for group size and directly measures selective amplification.
+Distributional Sharpness (Entropy):
+To formalize whether semantic alignment is concentrated or distributed, we compute the entropy over the normalized top- logits:
+Lower entropy indicates sharper specialization (typical of autoregressive models), while higher entropy indicates more distributed alignment (typical of masked diffusion models).
+C.2 Architectural Depth Metric
+Center of Gravity (CoG):
+To provide an objective, scalar measure of the “Mechanism Shift” independent of graph layout and visual density, we calculate the Center of Gravity (CoG).
+The CoG represents the attribution-weighted average layer index of the discovered circuit, identifying where the core computations are localized:
+where is the layer index and is the sum of EAP-IG attribution scores for all components in layer .
+The tables below provide the empirical results derived from the metrics defined above.
+We observe that autoregressive models exhibit larger logit gaps (e.g., Qwen: 2.12; LLaMA: 0.70) than their masked diffusion counterparts (DiffuLLaMA: 0.49), indicating a stronger concentration of probability mass on a single dominant token.
+Furthermore, DiffuLLaMA exhibits higher entropy (2.01) than autoregressive models (LLaMA: 1.81; Qwen: 0.97), indicating more distributed semantic alignment.
+Together, these metrics demonstrate the differences in specialization, selective amplification, and architectural depth between ARMs and MDMs.
+| Model | Name Token Proportion (Top-1) | Median LME | Logit Gap | Entropy |
+| Dream | 19.7% | 0.31 | – | – |
+| DiffuLLaMA | 25.2% | 0.05 | 0.49 | 2.01 |
+| Qwen | 6.0% | 0.90 | 2.12 | 0.97 |
+| LLaMA | 13.4% | 0.89 | 0.70 | 1.81 |
+| Task | Model | CoG (Layer Index) | Relative CoG (0–1) |
+| IOI | Qwen2.5-7B | 17.528 | 0.548 |
+| Dream-Base-7B | 20.432 | 0.638 | |
+| Countdown | Qwen2.5-7B | 16.582 | 0.518 |
+| Dream-Base-7B | 4.856 | 0.152 |